Object is thrown by 50 meter per second at an angle 30 degrees above horizontally find the time required to hit the ground


1
Expert's answer
2021-05-30T13:31:48-0400

"v_y=v_0\sin30°-gt\to 0=v_0\sin30°-gt"


"t=\frac{v_0\sin30°}{g}"


"t_{total}=2t=2\cdot\frac{v_0\sin30°}{g}=2\cdot\frac{50\sin30°}{9.8}=5.1\ (s)" . Answer